OpenStudy (anonymous):

Simplify, using order of operations: 6 + [3 2 + (4 + 9 3)]

OpenStudy (anonymous):

Simplifying 6 + (32 + (4 + 93)) = 0 Combine like terms: 4 + 93 = 97 6 + (32 + (97)) = 0 6 + (32 + 97) = 0 Combine like terms: 32 + 97 = 129 6 + (129) = 0 6 + 129 = 0 Combine like terms: 6 + 129 = 135 135 = 0 Solving 135 = 0 Couldn't find a variable to solve for. This equation is invalid, the left and right sides are not equal, therefore there is no solution.